the width of a rectangle is 5 meters less than its length. the area is 84 square meters. find the dimensions of the rectangle.
Answer by richwmiller(17219)

This problem could be in word problems geometry or quadratic equations.
Area=L*W
W=L-5
substitute L-5 for W
84=L*(L-5)
84=L^2-5L
L^2-5L-84=0
we ant factors of 84 with a difference of 5
1,84
2,42=6*7=3*2*7
3,14
4,21
6,14
7,12
7 and 12 will do
We want the larger number to be negative
(L-12)(L+7)=0
L=12 since a rectangle can't have a negative size of 7
W=L-5=7